OpenStudy (anonymous):

16. If the temperature of 50.0 L of a gas at 40OC falls to 30OC, what is the new volume of the gas if the pressure is constant? a. 45.0 L b. 52.0 L c. 55.0 L d. 48.4 L

OpenStudy (frostbite):

Okay. We work with a gas, lets assume it is a ideal gas we can then write the ideal gas equation as\[\Large pV=nRT\] Now in this gas we do not let the gas out so the amount of molecules \(n\) is constant. The ideal gas constant is as the name suggest also constant. For a condition change from one state to another we can then write the ideal gas equation as: \[\Large \frac{ p _{0}V _{0} }{ T _{0} }=\frac{ p _{1}V _{1} }{ T _{1} }\] It turns out that the pressure remains constant \(p_{0}= p_{1} \), so we can cancel the pressure: \[\Large \frac{ V _{0} }{ T _{0} }=\frac{ V _{1} }{ T _{1} }\] Set into the equation and evaluate the volume.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

so plug in the numbers

OpenStudy (anonymous):

I think there isn't a right answer

OpenStudy (anonymous):

so wut would it be then

OpenStudy (frostbite):

I think you should stick to my method: Isolate V_2 \[\Large V_{1}=\frac{ V _{0}T _{1} }{ T _{0} }\] Convert the Celsius to Kelvin. 30O C = 573 K 400 C = 673 K Evaluation: http://bit.ly/Oywnow